### v2

# Rethinking Early-Fusion Strategies for Improved Multispectral Object Detection

Abstract
Most recent multispectral object detectors employ a two-branch structure to extract features from RGB and thermal images. While the two-branch structure achieves better performance than a single-branch structure, it overlooks inference efficiency. This conflict is increasingly aggressive, as recent works solely pursue higher performance rather than both performance and efficiency. In this paper, we address this issue by improving the performance of efficient single-branch structures. We revisit the reasons causing the performance gap between these structures. For the first time, we reveal the information interference problem in the naive early-fusion strategy adopted by previous single-branch structures. Besides, we find that the domain gap between multispectral images, and weak feature representation of the single-branch structure are also key obstacles for performance. Focusing on these three problems, we propose corresponding solutions, including a novel shape-priority early-fusion strategy, a weakly supervised learning method, and a core knowledge distillation technique. Experiments demonstrate that single-branch networks equipped with these three contributions achieve significant performance enhancements while retaining high efficiency. Our code will be available at https://github.com/XueZ-phd/Efficient-RGB-T-Early-Fusion-Detection.
Index Terms:
Multispectral object detection; feature fusion; weakly supervised learning; knowledge distillation
I Introduction
Multispectral object detection has been widely studied, since multispectral images can provide complementary information to achieve consistent detection in various lighting conditions [3, 4, 5, 6, 7, 8, 9, 10]. This complementarity is illustrated in Fig. 1 (a) and (b). Given the multispectral inputs, modern multispectral detectors develop three fusion strategies: early-fusion, medium-fusion, and late-fusion, shown in Fig. 1 (c) - (e). The medium and late-fusion strategies often achieve superior performance compared to early-fusion [11, 12, 13, 5, 4, 14]. However, they use a two-branch structure, making model deployment on edge devices expensive. In contrast, the early-fusion strategy adopts a simple single-branch structure, facilitating deployment on edge devices. Nevertheless, its performance is low, and there are few works to address this problem, resulting in an increasing gap between high performance and high efficiency.
The main motivation for this work is to resolve the conflict between detection performance and inference efficiency. To this end, we focus on improving the performance of the early-fusion strategy while maintaining its high computational efficiency. We first conduct pilot studies and observe that a plain early-fusion strategy cannot consistently obtain improved performances compared to single-modality inputs. Based on this observation, we rethink the early-fusion strategy and summarize three key obstacles: 1) the information interference problem when simply concatenating the multispectral images, 2) the domain gap existing in thermal and RGB images, and 3) the weak feature representation of the single-branch structure. Focusing on these obstacles, we propose corresponding solutions.
- Information interference problem refers to the potential suppression of important information in one modality by another. In the plain early-fusion strategy, previous works [15] typically feed concatenated multispectral images into a convolution layer and generate a fused feature. The convolution layer generally has a small receptive field. Therefore, based on limited contexts, this approach is hard to determine which modality information is important. We address this issue by first recognizing that object shapes are agnostic to visible and infrared wavelengths and devise a module to fuse multispectral images based on object shape saliency, named the shape-priority early-fusion (ShaPE) module.
- Domain gap between RGB and thermal images is usually neglected in previous works. They generally adopt an RGB pre-trained backbone network to extract features from both RGB and thermal images [13, 5]. However, the domain gap may cause the representation distribution shift. This issue is also recognized in the work [16] on an RGB-D task. Different from previous works, we introduce a weakly supervised learning method to address this issue. Within this method, the backbone network jointly uses RGB and thermal images to learn the representation of CLIP [17], since CLIP has demonstrated promising zero-shot generalizability in bridging the domain gap [18]. Additionally, we introduce a segmentation auxiliary branch. Our method allows the backbone network to reduce representation shifts and improve semantic localization ability.
- Weak feature representation problem results from the early-fusion strategy employing a single-branch structure. This structure has fewer parameters and simpler fusion modules compared to medium and late-fusion strategies. We address this issue by introducing the knowledge distillation (KD) technique [19]. In KD, a key problem is how to align the feature dimensions between teacher and student models. Previous works generally introduce a convolution layer for the student model to learn all knowledge from the teacher model [20, 21]. However, we show that not all information in teacher model is helpful for downstream tasks. Therefore, we introduce core knowledge distillation (CoreKD) to transfer the most crucial knowledge for specific downstream tasks, resembling the human learning process where the teacher highlights key knowledge for quick understanding and absorption by the students.
Experimental results validate that our efficient multispectral early-fusion (EME) detector achieves a significant performance improvement without considerably increasing the number of parameters, as shown in Fig. 1 (f). Besides, our EME outperforms the previous state-of-the-art approaches. In summary, our contributions are threefold:
We systematically analyze the causes of the performance gap between single-branch and two-branch structures. Unlike previous works, we identify and summarize three key obstacles limiting the single-branch early-fusion strategy: information interference, domain gap, and weak feature representation. Notably, information interference between multispectral images is revealed for the first time in this work.
For each obstacle, we propose the corresponding solution: we develop 1) a ShaPE module to address the information interference issue, 2) a weakly supervised learning method to reduce domain gap and improve semantic localization abilities, and 3) a CoreKD to enhance the feature representation of single-branch networks.
Extensive experiments validate that the early-fusion strategy, equipped with our ShaPE module, weakly supervised learning, and CoreKD technique, demonstrates significant improvement. These three modules benefit various common detectors, such as YOLOv5 [2], RetinaNet [22], and GFL [23]. Importantly, only the ShaPE module is retained during the inference phase. Consequently, our method achieves both high performance and efficiency.
II Related Work
In this section, we offer a brief overview of multispectral object detection and introduce related works in weakly supervised learning and knowledge distillation.
II-A Multispectral Object Detection
Multi-source information fusion [24, 25, 26, 27] has exhibited promising application potential in computer vision tasks. In this work, we focus on the multispectral object detection task that uses RGB and thermal image pairs to detect objects. According to fusion strategies, multispectral object detection can be classified into three categories: early-fusion, medium-fusion and late-fusion strategies. Previous works [11], [12] and [28] confirm that both medium-fusion and late-fusion strategies outperform the early-fusion strategy.
However, both the medium and late fusion strategies adopt a two-branch structure that limits their use on resource-limited edge devices. Previous works notice this weakness and provide some solutions. For example, in [14], a model using the medium-fusion strategy is first trained as a teacher, and its knowledge is transferred to a student model. The student model only receives RGB images as inputs. Although it saves resources, it discards important complementary information from thermal images. The work [13] introduces a domain adaptation technique. It uses a medium-fusion model to guide single-branch model learning, which only receives thermal images as inputs and also discards complementary information from RGB images. To employ complementary information while saving computational resources, [29] transfers knowledge from a medium-fusion model to an early-fusion model. Nevertheless, it neglects information interference problem. Some works in the image fusion field [1, 30, 31] demonstrate that fused images can improve detectors, but the fusion process still introduces an additional computational burden.
Different from previous works, we identify the information interference problem in early-fusion strategies. By addressing this problem, we fully employ the complementary information in multispectral images, without significantly increasing computational burden.
II-B Weakly Supervised Learning in Object Detection
Weakly supervised learning has received much attention in object localization and detection, as comprehensively surveyed in [32]. Recent works in the multispectral object detection adopt this technique. Based on the weak annotations they utilize, we can coarsely divide them into image- and box-level weakly supervised learning approaches.
In image-level weakly supervised learning approaches, previous works mainly employ the illumination condition of RGB images as weighting factors to determine the modality importance [14, 5, 33, 13]. In box-level approaches, previous works [15, 34] mainly employ the bounding-box annotations to generate masks. They use these masks to construct spatial attention mechanisms, highlighting representations within target regions.
Different from previous works, we use weakly supervised learning to address the domain gap problem in RGB and thermal images. We employ image-level labels to construct a multi-label classification auxiliary task. This task can fully exploit the complementary information in multispectral images, instead of solely using information from one modality. Along with the powerful CLIP model [17] and box-level weak labels, our method can reduce the domain gap and obtain precise semantic localization abilities.
II-C Knowledge Distillation
Knowledge distillation is first introduced in [19]. It aims to improve a lightweight student model by learning knowledge from a high-capacity teacher model. According to distillation approaches, this technique can be roughly divided into two groups: logit distillation [19] and feature distillation [20]. The former let a student model learn the logit of a teacher model, while the latter let a student model learn the feature of a teacher model. These distillation approaches are also applied to object detection [35, 21]. Recently, some works in multispectral object detection also employ the knowledge distillation technique [29, 14]. In the distillation process, they generally introduce a projection layer to align the teacher and student feature channel number. The purpose of this approach is to learn all representations in the teacher model.
Different from previous works, we first confirm that not all information in teacher features is beneficial to downstream task including classification and regression. Based on this, we propose a core knowledge distillation technique to transfer the most important features for the downstream tasks to the student model.
III Method
Fig. 2 illustrates the overview of our method, where the training process and the inference process are marked in green and blue, respectively. We adopt a single-branch structure as the baseline model considering its low memory cost. To boost its performance, we develop three key modules: shape-priority early-fusion (ShaPE), weakly supervised auxiliary learning, and core knowledge distillation (CoreKD). Note that only the training process requires weakly supervised auxiliary learning and CoreKD, and both are removed during the inference phase. Consequently, our method adds only the ShaPE module to the early-fusion single-branch structure during the inference phase. In the following sections, we describe the ShaPE module in Section III-A, the weakly supervised auxiliary learning method in Section III-B, and CoreKD in Section III-C.
III-A Shape-Priority Early-Fusion Module
Observation. Given a pair of RGB-T images, the plain early-fusion strategy concatenates them in the channel dimension and then feeds them into a detector. With the plain strategy, we conduct pilot studies on the M3FD [1] dataset. We first train three commonly used one-stage detectors: RetinaNet [22], GFL [23] and YOLOv5 [2]. Then, we compute the mean values and standard deviations of their detection results and illustrate the computed results in Fig. 3. Besides, we also train these detectors using single-modality images as input for comparisons. We have the following two observations. First, the plain early-fusion strategy cannot achieve consistent improvement compared with single-modality input. Second, for objects that require color to identify, such as ‘Traffic Light’, the plain early-fusion strategy yields worse results than the RGB input.
Motivation. We attribute the above phenomena to the convolutional inductive bias, namely, local connectivity and weight sharing. The process of 2D convolution involves two steps: (1) sampling across the concatenated RGB-T images using a regular grid ; (2) summing the sampled values with weighting factor . The grid determines both the receptive field size and dilation. For example,
defines a 77 kernel with dilation 1. For each position on an out feature map , we have
where enumerates the positions in .
This process indicates that the plain early-fusion strategy is a pixel-level weighting method, with weights learned from data. However, the limited receptive field of pixel-level weighting methods makes the weights difficult to determine which modality is important. This weakness may result in valuable information from one modality being suppressed by another. As an example, Fig. 4 (c) depicts the feature map generated from the RGB-T images of Fig. 4 (a) and (b) using the plain early-fusion strategy. It is observed from the close-up that the ‘Traffic Light’ in the fused feature map doesn’t preserve the significant information of the RGB image.
The straightforward solutions to this weakness are: (1) enlarging the receptive field by using a larger kernel or more convolutional layers so that the model can judge the modality importance based on a broader range of contexts, or (2) increasing the number of convolutional kernels so that the model can learn more representations. However, these solutions increase memory costs and computational burden, making them unfriendly to edge devices.
ShaPE Module. We realize that shape is an inherent attribute of an object. Any visible objects in RGB and thermal images have consistent shapes. Thus, we consider the salience of shape as a modifying factor to adaptively determine the modality importance, and design the shape-priority early-fusion (ShaPE) module. In the ShaPE module, the RGB and thermal images are modified by self-gating masks. In this context, Eq. (1) becomes:
where and denote the self-gating masks of RGB and thermal images, respectively.
In the following, we describe the generation process of self-gating masks and . Since our ShaPE module focuses on the shapes of objects and structural contributions of different modalities to the fused features, we employ the gradients and structural similarities in our method. For easy understanding, we visualize some important intermediate results in Fig. 4. Given the RGB-T images as shown in Fig. 4 (a) and (b), we compute their gradients
as shown in Fig. 4 (d) and (e). We then generate the union gradient as the reference using
We further use max-pooling within a 33 neighborhood to boost the reference gradient, which is written as
as shown in Fig. 4 (f).
To determine the structural contributions of each modality to the fused features, we compute the structural similarities between single-modality gradient images {, } and the reference gradient image . Inspired by [36], for each patch , we compute three fundamental properties: the means {}, the standard deviations {}, and the covariances {, } between the single-modality gradient images and the reference gradient images. In this context, we generate the self-gating masks:
where and are used to prevent instability. is the dynamic range of the gradient images, , and .
III-B Weakly Supervised Learning Method
In RGB-T object detection, an unneglectable issue is the lack of pre-trained backbone networks on large-scale RGB-T datasets. This is because there are few large-scale datasets like ImageNet [37] and COCO [38] in RGB-T image recognition fields. Previous works generally use backbone networks pre-trained on ImageNet. However, the domain gap between thermal and RGB images would cause representation distribution shifts, as illustrated in Fig. 5 (a) and (b). This is because the backbone network is trained solely on RGB images, but is applied to thermal images.
To handle this issue, we turn to the powerful Contrastive Language-Image Pre-training (CLIP) [17] model. It has been confirmed that CLIP can bridge domain gaps [18, 40, 41, 42, 43], since it is trained using a huge number of (image, text) pairs. In this context, we feed both RGB and thermal images into the backbone network, and let it learn the representation generated by the CLIP model. Specifically, we first present a CLIP-driven image-level weakly supervised learning method. This method enables the network to recognize the classes of objects in a pair of RGB-T images while locating their coarse regions. For fine-grained localization, we then introduce a box-level weakly supervised learning method. Fig. 6 illustrates the architecture of weakly supervised learning method.
CLIP-Driven Image-Level Weak Supervision. To learn the CLIP model’s knowledge, we construct the image-level weak supervision method. Based on three considerations, we adopt the multi-label classification task as the image-level weak supervision: (1) the CLIP model can be viewed as a classifier, (2) this auxiliary task can fully use the complementary information in the RGB-T images, and (3) by summarizing all classes and removing duplicates in an image, we can easily construct the ground-truth multi-label targets based on detection annotations.
Nevertheless, original CLIP model is only trained for recognizing a single object per image [17] and is not suitable for multi-label classification [44, 45]. To address this issue, we introduce a Divide-and-Aggregation CLIP (DA-CLIP) model. DA-CLIP first divides input images into multiple crops. Each crop is then fed into CLIP. All predictions of these crops are finally aggregated by a max-pooling operation on each class. Considering DA-CLIP may generate inaccurate predictions, we construct a learnable adapter, which consists of three fully-connected (FC) layers, to fine-tune the result of DA-CLIP. To prevent overfitting, we add a dropout layer in the adapter. We denote the predicted probability from the adapter as , where denotes the number of classes.
For the backbone network, we add an auxiliary classification head on its top. The head consists of a global average pooling (GAP) operation and one FC layer. We denote the predicted probability from the classification head as .
We adopt the mutual learning approach [46] to train the backbone network and the adapter simultaneously. In this approach, an important step is that one model generates soft targets for the other model using the softmax function. However, this approach cannot be directly applied to the multi-label classification problem, since it requires the sum of predicted probabilities to be one, which is rarely satisfied in multi-label classification. To address this issue, we draw inspiration from self-training KD [47] and construct the soft targets for the adapter and backbone network as
where denotes a ground-truth multi-label target, and denotes a balancing factor set to 0.1. In this context, we compute the binary cross-entropy (BCE) losses
To showcase the semantic localization effect of our CLIP-driven image-level weak supervision, we visualize the class activation map (CAM) of the backbone network in Fig. 8 (a). CAM is a useful tool for understanding which regions the network focuses on to predict a class. We can observe that the backbone network can coarsely localize regions of ‘Person’, ‘Car’, and ‘Traffic Light’ in the image.
Box-Level Weak Supervision. To precisely localize the semantic regions, we introduce box-level weak supervision. The ground-truth box-level target is generated by directly filling the area within an annotation box with its corresponding class index. In this context, we add an auxiliary segmentation head on top of the backbone network to predict the target. Denoting the ground-truth box-level target mask as , and the predicted mask as , we compute the BCE loss between them as
where denotes the number of elements in the mask.
We visualize attention maps of the backbone network for different classes, as shown in Fig. 8 (b). Using the box-level weak supervision, the backbone network can precisely localize the interest of objects, such as ’Car’. Nevertheless, it may miss some useful information in the image. Therefore, we combine the CLIP-driven image-level weak supervision and the box-level weak supervision. The results presented in Fig. 8 (c) show that our weakly supervised learning method can effectively allow the backbone network to localize the important semantic regions.
Effect Validation. When our weakly supervised learning method is employed, Fig. 5 (c) and (d) demonstrate that the domain gap between RGB and thermal features is reduced. This implies that the backbone network can extract information from RGB and thermal images without bias. To further illustrate this effect, we visualize the feature map generated by the ResNet-50 [48] in Fig. 8. The generation process of these feature maps is as follows: First, we resize all features of the ResNet-50 across four stages to the same resolution as the input images. Then, we aggregate these features along the channel dimension using , where represents the concatenated feature. , , and denote its depth, height, and width, respectively. denotes the element-wise production operation.
Fig. 8 (a) and (b) present the RGB and thermal images in one example scene. Fig. 8 (c) and (d) illustrate their corresponding feature maps. Fig. 8 (e) shows the RGB-T feature map without using our weakly supervised learning method. Fig. 8 (f) shows the feature map using our weakly supervised learning method. Observing Fig. 8 (e), we note that the ResNet-50 tends to acquire information primarily from the RGB image. In contrast, the feature map in Fig. 8 (f) demonstrates that our method enables the ResNet-50 to gather important information from both RGB and thermal images.
III-C Core Knowledge Distillation
Problem Description. To further improve the detection accuracy of the early-fusion strategy without increasing its computational cost, we introduce the knowledge distillation technique [19]. To achieve knowledge transfer, we instruct the student model to mimic intermediate features of teacher model. In this process, a primary obstacle the student model faces is the unequal number of feature channels as the teacher model. Previous works introduce convolution layers to align their feature channel numbers [20, 21], while neglecting whether the teacher’s knowledge is helpful to the student. To address this issue, we propose core knowledge distillation (CoreKD).
CoreKD Architecture. We use YOLOv5 [2] as an example and illustrate the knowledge distillation architecture in Fig. 9. In its architecture, we use the early-fusion single-branch structure as the student model and the medium-fusion two-branch structure as the teacher model. In the student model, a pair of RGB-T images is first concatenated, then fed into different network modules, and finally converted into predicted results. In the teacher model, the RGB and thermal images are respectively fed into different backbone networks. The generated multispectral features are fused in the feature space through concatenation and convolution operations. The fused features are then fed into the subsequent network modules and converted into predicted results. The predicted results of both the student and teacher models consist of bounding boxes and class-specific confidence scores.
CoreKD Formulation. Since we apply the same distillation techniques to different feature pyramid levels, we only describe the technique at one level and omit the subscript for simplicity. In the head modules of Fig. 9, we denote the input features of the student and teacher models as and , respectively. Feature distillation typically transfers the teacher’s knowledge to the student by minimizing the loss [20]
where denotes an adaptation layer used to match the channel dimensions between the student and teacher features. Previous works usually use a convolution layer as the adaptation layer [20, 21]. This approach aims to make learn all information in the teacher feature . However, they neglect whether all the information in is beneficial for downstream tasks, including classification and regression.
To address this problem, we revisit the structure of head module in the teacher model. As shown in Fig. 9, the official implementation of YOLOv5 uses a ‘’ layer to output the predicted results
where denotes the weighting factor in the teacher’s head module. According to the 2D convolution formulation in Eq. (1), we can infer that the weighting factor reflects the importance of a channel map in for the downstream feature. We visualize the histogram of in Fig. 10. It is evident that most of the values in approximate 0. This implies that only a few feature representations in are important for the downstream tasks. We call these important feature representations the core knowledge in teacher model.
To learn this core knowledge, we modify the feature loss Eq. (6) into
This modification ensures that and are projected into an identical space constructed by , and that the projected features are close to each other. Furthermore, to avoid introducing the adaption layer , we construct a core knowledge convolution (Core Knowledge Conv) operator by sampling the weighting factor . We denote the sampling process as . In the process, we first obtain the channel dimension of the student feature , then sample the top- values along the ‘in_channel’ axis from based on their absolute values. Finally, we obtain the sampled weighting factor . In this context, we rewrite the feature loss given in Eq. (7) as
where denotes the output of core knowledge convolution. When using this feature loss, we keep the weighting factor fixed and only compute the gradient with respect to the student feature .
Mathematical Foundation of CoreKD. We first explain the mathematical foundation of traditional feature distillation and analyze its weaknesses. Then, we introduce the mathematical foundation of our CoreKD. Finally, we compare the results of our CoreKD with those of the traditional one.
Traditional feature distillation uses a convolution layer to align feature channel numbers between student and teacher models. We denote the convolution layer as a function in the Eq. (6). Next, we denote the weighting factor of as . This function is used to convert an input feature into an output feature , i.e., . We denote the vector at an arbitrary spatial location of as , the ith row vector in the weighting factor as , and the corresponding value in as . The mathematical relation between , , and can be represented as
where the operator ‘’ indicates a dot product. The dot product computation can be viewed as the projection of the vector onto the vector , as shown in Fig. 11 (a). We can infer that the generation of is related to the and but has no relation to the teacher’s features. This implies that the traditional feature distillation merely focuses on enforcing the student to learn all information from the teacher without considering whether the teacher’s features are beneficial to downstream tasks.
On the contrary, our CoreKD uses the weighting factor of the teacher model to align feature channel numbers between the student and teacher models, as shown in Eq. (8). We denote the ith row vector of as and the vector at an arbitrary spatial location of as . Then we can write the ith loss value of as
This loss value calculation process is illustrated in Fig. 11 (b). From the above analyses, we have two key observations: 1) our CoreKD projects the student and teacher features into an identical space constructed by , and 2) our CoreKD does not enforce the student feature to be the same as the teacher feature but rather focuses on minimizing the projected distances. Since the values within the weighting factor reflect the importance of teacher features, our CoreKD enables the student model to learn beneficial information for downstream tasks from the teacher model.
Since the experimental results involve the introduction of both datasets and implementation details, we arrange the comparison results of our CoreKD with the traditional feature distillation in the experimental section. For details, please refer to Section IV-C.
III-D Loss Function
Our efficient multispectral early-fusion (EME) single-branch model is trained using all the losses described above. The total loss is
where and represent the classification and regression losses defined by a detector [22, 23, 2], respectively. is the summation of weakly supervised losses defined in Eq. (4) and Eq. (5):
IV Experiments
IV-A Experimental Setup
Datasets. Our experiments are conducted on the M3FD dataset [1] and FLIR dataset [39]. M3FD dataset contains 4200 pairs of RGB and thermal images. These image pairs are well aligned. The dataset contains 6 classes of objects: ‘Person’, ‘Car’, ‘Bus’, ‘Motorcycle’, ‘Traffic Light’, and ‘Truck’. Since this dataset doesn’t provide unified data splits, previous works have used a random splitting approach to determine the train and validation sets [1]. However, images in this dataset are sampled from video sequences, meaning that two adjacent frames may contain identical content. In this context, the random splitting approach results in information leakage between the train and validation sets. To address this problem, we first manually divide the dataset into 73 video segments based on different scenes. Then, we collect the first 70% of images in each video segment as the train set and the remaining images as the validation set. Finally, we obtain 2905 and 1295 pairs of RGB-T images in the train and validation sets, respectively. We name this data split ‘M3FD-zxSplit’ and release it to the public111https://github.com/XueZ-phd/Efficient-RGB-T-Early-Fusion-Detection. For the performance evaluation in Section IV-B, we use this data split. When comparing with state-of-the-art approaches in Section IV-D, we employ both ‘M3FD-zxSplit’ and random splitting. Our random splitting refers to randomly selecting 80% images as the train set and the remaining images as the validation set. FLIR dataset originally contains unaligned RGB-T image pairs. The work [49] develops a data-processing approach to align these images and obtain 7381, 1056, and 2111 image pairs in the train, validation, and test sets. This dataset contains 3 classes: ‘Person’, ‘Bicycle’ and ‘Car’.
Evaluation Metrics. We use the standard mean Average Precision (mAP) with IoU thresholds ranging from 0.5 to 0.95 across various object scales as metrics.
Inference Efficiency Evaluations. We assess the inference efficiency of our method (Python implementation) on the edge device NVIDIA AGX Orin with 64GB memory. We also evaluate the complexity of our method using FLOPs and the number of parameters. All results are presented in Tables I and II.
Implementation Details. We incorporate our three key modules into commonly-used one-stage detectors, including RetinaNet [22], GFL [23], and YOLOv5 [2]. For RetinaNet and GFL, we adopt the implementations in MMDetection toolbox [50]. For YOLOv5, we use its official implemtation [2].
In the early-fusion strategy based on RetinaNet [22] and GFL [23] detectors, we use ResNet-50 [48] as the backbone network. For a fair comparison, we use the same backbone network in the medium-fusion strategy. Notably, in the CoreKD technique, the teacher model utilizes ResNet-101 [48] as the backbone network. For both strategies, we train for 12 epochs using the SGD optimizer with a batch size of 4. The initial learning rate is set to 0.01 and is decayed by 0.1 at epochs 8 and 11. Random horizontal flipping is employed as a data augmentation technique.
For the early-fusion strategy based on YOLOv5, we use YOLOv5-small as the baseline detector, and we use YOLOv5-large to construct a teacher model in the CoreKD technique. For both strategies, we train for 36 epochs with a batch size of 16. We keep all other hyperparameters consistent with the official settings of the YOLOv5 repository [2].
To standardize data for RetinaNet and GFL detectors, we calculate the mean value and standard deviation of RGB and thermal images for the M3FD and FLIR datasets. All experiments use the 640 512 image resolution. For the M3FD dataset, we obtain meanrgb = [128.2, 129.3, 125.3], stdrgb = [49.1, 50.2, 53.5], meant = [84.1, 84.1, 84.1], and stdt = [50.6, 50.6, 50.6]. For the FLIR dataset, we obtain meanrgb = [149.4, 148.7, 141.7], stdrgb = [49.3, 52.8, 59.0], meant = [135.7, 135.7, 135.7], and stdt = [63.6, 63.6, 63.6]. For the YOLOv5 detector, we normalize both RGB and thermal images to the range of [0, 1] following its official implementations.
Detector | FLOPs (↓) | Parameter (↓) | Time (↓) | mAP (↑) | mAP50 (↑) | Person (↑) | Car (↑) | Bus (↑) | Motor (↑) | TrafficLight (↑) | Truck (↑) | |
RGB | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 31.03±0.09 | 51.30±0.16 | 44.57±0.12 | 74.87±0.26 | 57.80±0.24 | 44.30±0.22 | 36.63±0.58 | 49.70±0.16 |
Thermal | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 29.27±0.05 | 46.97±0.09 | 59.17±0.34 | 71.17±0.29 | 54.17±0.69 | 35.90±0.22 | 10.43±0.09 | 50.83±0.21 |
RGB-T Medium Fusion | RetinaNet-Res50 | 94.611G | 47.582M | 0.170s | 33.43±0.05 | 53.63±0.05 | 60.10±0.08 | 77.27±0.05 | 61.63±0.05 | 45.43±0.09 | 25.67±0.12 | 51.80±0.00 |
Baseline: RGB-T Early Fusion | RetinaNet-Res50 | 62.164G | 36.434M | 0.110s | 32.03±0.05 | 50.70±0.16 | 58.93±0.26 | 75.37±0.09 | 58.97±0.39 | 39.97±0.21 | 21.67±0.47 | 49.33±0.05 |
+ ShaPE | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 32.80±0.08 | 51.97±0.24 | 60.20±0.59 | 77.10±0.08 | 58.83±0.33 | 39.77±0.58 | 24.57±0.54 | 51.33±0.17 |
+ ShaPE + WeakSup. | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 33.53±0.05 | 52.90±0.16 | 59.10±0.71 | 77.07±0.41 | 62.00±0.78 | 40.97±1.67 | 24.03±0.85 | 54.30±0.71 |
+ ShaPE + WeakSup. + CoreKD | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 33.53±0.17 | 53.23±0.09 | 61.47±0.49 | 76.40±0.08 | 59.20±0.08 | 43.10±0.22 | 25.97±0.12 | 53.17±0.19 |
RGB | GFL-Res50 | 61.392G | 32.270M | 0.110s | 32.23±0.12 | 53.10±0.16 | 48.67±0.12 | 77.43±0.26 | 60.27±0.60 | 43.50±0.16 | 39.07±0.94 | 49.63±0.39 |
Thermal | GFL-Res50 | 61.392G | 32.270M | 0.110s | 29.50±0.22 | 48.27±0.42 | 64.27±0.34 | 73.73±0.12 | 52.50±1.99 | 36.50±0.0 | 15.10±0.14 | 47.37±0.25 |
RGB-T Medium Fusion | GFL-Res50 | 94.110G | 43.419M | 0.172s | 34.17±0.33 | 54.47±0.60 | 65.37±0.12 | 79.83±0.05 | 61.20±1.00 | 37.00±1.79 | 34.80±0.57 | 48.73±0.77 |
Baseline: RGB-T Early Fusion | GFL-Res50 | 61.663G | 32.271M | 0.114s | 33.50±0.28 | 52.77±0.25 | 64.03±0.12 | 78.20±0.16 | 53.00±1.31 | 39.27±0.26 | 30.47±0.56 | 51.57±0.54 |
+ ShaPE | GFL-Res50 | 61.718G | 32.271M | 0.151s | 35.17±0.17 | 55.50±0.22 | 65.80±0.24 | 79.10±0.08 | 62.33±1.24 | 41.33±0.42 | 30.80±0.42 | 53.67±0.41 |
+ ShaPE + WeakSup. | GFL-Res50 | 61.718G | 32.271M | 0.151s | 35.23±0.53 | 55.97±0.24 | 65.73±0.49 | 79.57±0.17 | 60.50±4.02 | 43.10±1.06 | 33.73±2.01 | 53.10±1.94 |
+ ShaPE + WeakSup. + CoreKD | GFL-Res50 | 61.718G | 32.271M | 0.151s | 37.03±0.09 | 57.70±0.08 | 68.43±0.25 | 81.23±0.09 | 63.37±0.33 | 43.90±0.86 | 35.77±0.12 | 53.53±0.29 |
Detector | FLOPs (↓) | Parameter (↓) | Time (↓) | mAP (↑) | mAP50 (↑) | Person (↑) | Bicycle (↑) | Car (↑) | |
RGB | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 28.10±0.0 | 59.47±0.12 | 44.93±0.39 | 55.70±0.08 | 77.80±0.08 |
Thermal | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 35.53±0.05 | 70.93±0.05 | 62.17±0.17 | 66.37±0.09 | 84.27±0.05 |
RGB-T Medium Fusion | RetinaNet-Res50 | 94.611G | 47.582M | 0.170s | 38.50±0.08 | 71.57±0.05 | 61.93±0.29 | 67.60±0.14 | 85.17±0.09 |
Baseline: RGB-T Early Fusion | RetinaNet-Res50 | 62.164G | 36.434M | 0.110s | 37.47±0.05 | 69.57±0.05 | 60.70±0.22 | 63.77±0.09 | 84.37±0.05 |
+ ShaPE | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.70±0.14 | 71.60±0.22 | 61.40±0.54 | 68.60±0.36 | 84.70±0.22 |
+ ShaPE + WeakSup. | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.80±0.29 | 72.07±0.37 | 62.50±0.96 | 68.77±0.97 | 85.03±0.21 |
+ ShaPE + WeakSup. + CoreKD | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.83±0.17 | 72.23±0.31 | 62.27±0.26 | 69.23±0.66 | 85.10±0.22 |
RGB | GFL-Res50 | 61.392G | 32.270M | 0.110s | 31.73±0.12 | 63.77±0.05 | 51.70±0.08 | 57.87±0.25 | 81.77±0.05 |
Thermal | GFL-Res50 | 61.392G | 32.270M | 0.110s | 42.40±0.14 | 75.07±0.05 | 69.80±0.16 | 68.47±0.24 | 86.93±0.05 |
RGB-T Medium Fusion | GFL-Res50 | 94.110G | 43.419M | 0.172s | 42.60±0.08 | 76.07±0.21 | 70.07±0.19 | 70.57±0.41 | 87.63±0.05 |
Baseline: RGB-T Early Fusion | GFL-Res50 | 61.663G | 32.271M | 0.114s | 41.90±0.22 | 74.77±0.17 | 69.70±0.33 | 67.73±0.26 | 87.00±0.00 |
+ ShaPE | GFL-Res50 | 61.718G | 32.271M | 0.151s | 42.40±0.16 | 75.77±0.17 | 69.97±0.21 | 70.13±0.37 | 87.23±0.09 |
+ ShaPE + WeakSup. | GFL-Res50 | 61.718G | 32.271M | 0.151s | 42.93±0.24 | 76.90±0.22 | 71.30±0.14 | 71.40±0.50 | 87.97±0.09 |
+ ShaPE + WeakSup. + CoreKD | GFL-Res50 | 61.718G | 32.271M | 0.151s | 44.00±0.00 | 78.17±0.05 | 73.03±0.05 | 72.63±0.17 | 88.80±0.00 |
IV-B Performance Evaluation of Proposed Modules
Table I and Table II present the performance of our method on the M3FD [1] and FLIR [39] datasets. Key observations include: (1) the medium-fusion strategy adds more parameters and FLOPs compared to the early-fusion strategy; (2) the medium-fusion strategy achieves better performance compared to single-modality inputs, whereas the plain early-fusion strategy does not consistently improve performance; (3) our EME method, incorporating the ShaPE module, weakly supervised learning, and CoreKD techniques into the plain early-fusion strategy, achieves significant performance improvement without significantly increasing parameters and FLOPs; (4) the inference time of our EME method is longer than that of the baseline method, since the structural similarity computation process has not been optimized when calculating the self-gating mask; (5) our EME method can outperform the medium-fusion strategy in both performance and efficiency to some extent; (6) Both architectures: “Baseline + ShaPE + WeakSup.” and “Baseline + ShaPE + WeakSup. + CoreKD” have the same FLOPs, parameters, and inference time as “Baseline + ShaPE”. This is because both the weakly supervised learning method and CoreKD are removed in the inference phase, while only the ShaPE module is retained.
Fig. 13 and Fig. 13 present visualization results for two example scenes from M3FD [1] and FLIR [39] datasets, respectively. We observe that false positives or false negatives in the single-modality results may affect the plain early-fusion strategy. For instance, the person missed in Fig. 13 (e) is also absent in Fig. 13 (g), despite being detected in Fig. 13 (f). Moreover, false positives in Fig. 13 (f) affect the detection results of plain early-fusion, as shown in Fig. 13 (g). These phenomena confirm that the problem of information interference is a key obstacle to performance in the plain early-fusion strategy. Clearly, our EME effectively alleviates this problem.
IV-C Performance Evaluation of Variants
Feature Distillation. We compare the results of traditional feature distillation with those of our CoreKD. Table III and Table IV present comparison results on the M3FD and FLIR datasets, respectively. For comprehensive comparisons, we adopt RetinaNet-Res50 and GFL-Res50 as baseline detectors in these two tables. From the results, we can observe that our CoreKD consistently achieves superior performance compared to traditional feature distillation. For example, our CoreKD (72.23%) obtains a 1.26% mAP50 absolute gain over the traditional one (70.97%) when using RetinaNet-Res50 on the FLIR dataset.
Method | Detector | mAP (↑) | mAP50 (↑) | ||
RetinaNet-Res50 | 32.03±0.05 | 50.70±0.16 | ||
RetinaNet-Res50 | 32.17±0.12 | 52.43±0.05 | ||
RetinaNet-Res50 | 33.53±0.17 | 53.23±0.09 | ||
GFL-Res50 | 33.50±0.28 | 52.77±0.25 | ||
GFL-Res50 | 35.83±0.17 | 57.07±0.09 | ||
GFL-Res50 | 37.03±0.09 | 57.70±0.08 |
Method | Detector | mAP (↑) | mAP50 (↑) | ||
RetinaNet-Res50 | 37.47±0.05 | 69.57±0.05 | ||
RetinaNet-Res50 | 38.10±0.14 | 70.97±0.12 | ||
RetinaNet-Res50 | 38.83±0.17 | 72.23±0.31 | ||
GFL-Res50 | 41.90±0.22 | 74.77±0.17 | ||
GFL-Res50 | 43.63±0.09 | 77.80±0.08 | ||
GFL-Res50 | 44.00±0.00 | 78.17±0.05 |
Backbone Network. We evaluate our EME method using ResNet-101 [48] as the backbone network on the M3FD and FLIR datasets, and present the results in Table V. We observe that detectors using ResNet-101 consistently achieve better performance than those using ResNet-50. For example, RetinaNet with ResNet-101 (74.87%) obtains a 2.64% mAP50 absolute gain over ResNet-50 (72.23%).
Datasets | Detector | Backbone | FLOPs | mAP (↑) | mAP50 (↑) |
M3FD | RetinaNet | ResNet-50 | 62.218G | 33.53±0.17 | 53.23±0.09 |
RetinaNet | ResNet-101 | 65.392G | 34.23±0.12 | 54.63±0.45 | |
GFL | ResNet-50 | 61.718G | 37.03±0.09 | 57.70±0.08 | |
GFL | ResNet-101 | 64.892G | 37.37±0.12 | 58.60±0.08 | |
FLIR | RetinaNet | ResNet-50 | 62.218G | 38.83±0.17 | 72.23±0.31 |
RetinaNet | ResNet-101 | 65.392G | 40.67±0.05 | 74.87±0.34 | |
GFL | ResNet-50 | 61.718G | 44.00±0.00 | 78.17±0.05 | |
GFL | ResNet-101 | 64.892G | 44.47±0.05 | 79.57±0.05 |
(a) Dataset Splitting Method: Random Splitting | |||||||||||||
Thermal [2] | RGB [2] | AUIF [51] | CDDF [30] | DDcGAN [52] | DIVF [31] | DenseF [53] | PSF [54] | RFN [55] | SeAF [56] | TarDAL [1] | U2F [57] | EME (Ours) | |
mAP | 49.10 | 52.40 | 53.30 | 53.00 | 52.20 | 52.70 | 53.40 | 53.10 | 53.50 | 53.10 | 52.50 | 53.40 | 54.00±0.28 |
mAP50 | 77.30 | 81.90 | 81.90 | 80.90 | 81.60 | 81.50 | 81.70 | 82.00 | 81.70 | 82.20 | 81.00 | 81.90 | 82.90±0.37 |
Person | 79.30 | 68.40 | 76.70 | 76.30 | 73.60 | 74.50 | 76.50 | 76.70 | 75.30 | 77.00 | 79.10 | 77.00 | 79.53±0.26 |
Car | 87.90 | 90.80 | 91.00 | 91.00 | 90.70 | 91.10 | 91.40 | 90.80 | 91.00 | 91.10 | 90.50 | 91.20 | 91.90±0.29 |
Bus | 87.20 | 92.20 | 90.00 | 90.10 | 90.70 | 91.60 | 89.40 | 90.10 | 89.40 | 91.20 | 89.40 | 90.70 | 89.80±0.45 |
Motor | 70.00 | 74.00 | 72.60 | 69.20 | 74.80 | 73.50 | 72.80 | 73.30 | 73.30 | 72.20 | 70.30 | 71.30 | 74.87±0.95 |
TrafficLight | 55.90 | 80.30 | 77.40 | 75.40 | 76.90 | 74.80 | 77.20 | 78.20 | 77.40 | 77.60 | 72.70 | 77.70 | 77.40±1.13 |
Truck | 83.40 | 85.70 | 83.70 | 83.10 | 82.90 | 83.40 | 82.90 | 82.90 | 83.90 | 84.10 | 84.00 | 83.60 | 84.00±0.93 |
(b) Dataset Splitting Method: M3FD-zxSplit | |||||||||||||
Thermal [2] | RGB [2] | AUIF [51] | CDDF [30] | DDcGAN [52] | DIVF [31] | DenseF [53] | PSF [54] | RFN [55] | SeAF [56] | TarDAL [1] | U2F [57] | EME (Ours) | |
mAP | 34.90 | 36.10 | 38.30 | 38.60 | 37.10 | 37.10 | 38.90 | 38.00 | 38.20 | 38.90 | 39.10 | 38.70 | 41.10±0.29 |
mAP50 | 57.20 | 60.20 | 62.00 | 61.90 | 61.00 | 60.80 | 62.40 | 61.10 | 61.30 | 62.20 | 61.90 | 61.90 | 66.23±0.40 |
Person | 74.60 | 55.90 | 72.20 | 71.90 | 67.30 | 67.60 | 72.30 | 71.70 | 70.50 | 72.50 | 75.50 | 72.40 | 77.23±0.26 |
Car | 80.20 | 84.80 | 85.50 | 85.60 | 84.90 | 85.20 | 85.90 | 85.50 | 85.80 | 85.50 | 85.00 | 85.50 | 87.13±0.19 |
Bus | 58.30 | 65.70 | 58.60 | 61.80 | 61.60 | 59.80 | 61.40 | 58.30 | 61.30 | 61.50 | 60.90 | 60.10 | 62.33±1.96 |
Motor | 48.00 | 45.10 | 49.10 | 47.60 | 49.00 | 48.70 | 49.60 | 45.80 | 44.60 | 47.50 | 46.80 | 50.80 | 55.33±0.21 |
TrafficLight | 27.30 | 56.80 | 49.80 | 48.70 | 49.10 | 51.20 | 48.60 | 50.90 | 49.70 | 50.80 | 46.90 | 48.00 | 55.33±0.37 |
Truck | 54.80 | 52.70 | 56.70 | 55.50 | 53.80 | 52.60 | 56.80 | 54.70 | 55.80 | 55.70 | 56.70 | 54.70 | 60.10±0.16 |
CBF [58] | MCG [59] | MUN [59] | ODS [60] | CFR [61] | GAFF [15] | BU [62] | SMPD [63] | ThDe [64] | MSAT [65] | CSAA [66] | MFPT [67] | ProbEn3 [28] | EME (Ours) | |
mAP50 | 67.20 | 61.40 | 61.54 | 69.62 | 72.39 | 72.90 | 73.20 | 73.58 | 74.60 | 76.20 | 79.20 | 80.00 | 83.76 | 84.63±0.12 |
Bicycle | 60.50 | 50.26 | 49.43 | 55.53 | 55.77 | - | 57.40 | 56.20 | 60.04 | - | - | 67.70 | 73.49 | 79.73±0.09 |
Car | 83.60 | 70.63 | 70.72 | 82.33 | 84.91 | - | 86.50 | 85.80 | 85.52 | - | - | 89.00 | 90.14 | 92.67±0.12 |
Person | 57.60 | 63.31 | 64.47 | 71.01 | 74.49 | - | 75.60 | 78.74 | 78.24 | - | - | 83.20 | 87.65 | 81.27±0.09 |
IV-D Comparison with the State-of-The-Art Approaches
We use the one-stage YOLOv5 [2] detector as the baseline, and incorporate our proposed modules to construct the effective multispectral early-fusion (EME) model. Table VI and Table VII compares our EME and previous state-of-the-art approaches on M3FD [1] and FLIR [39] datasets.
In Table VI, we compare our EME with 10 state-of-the-art image-fusion-based object detection approaches [51, 30, 52, 31, 53, 54, 55, 56, 1, 57]. We first generate fused images based on their official implementations, and then train YOLOv5 [2] using these fused images with the same training settings. The results show that our EME achieves state-of-the-art performance. We observe that the results in Table VI (a) are obviously better than those in Table VI (b). This demonstrates that random splitting causes information leakage and makes it difficult to improve performance. Fig. 14 presents an example scene for visualization. Compared to other approaches, a weakness of our EME detector is that it doesn’t generate a fused image for direct visualization. This is because our method focuses on detection rather than image fusion. We will address this issue in future work.
In Table VII, we compare our EME with 13 multispectral object detection approaches. These approaches include (1) medium-fusion strategies, such as CBF [58], MCG [59], MUN [59], CFR [61], GAFF [15], SMPD [63], MSAT [65], CSAA [66], and MFPT [67]; (2) domain adaptation and single-modality detection approaches, such as ODS [60], BU [62], and ThDe [64]; and (3) late-fusion strategy [28]. The results show that our EME also achieves state-of-the-art performance on the FLIR dataset [39].
IV-E Comparison of Inference Efficiency
We compare the inference efficiency of our EME method with previous state-of-the-art approaches on an edge device: the NVIDIA AGX Orin with 64GB of memory. We select open-source approaches for comparison and adopt YOLOv5-small as the baseline detector for all methods.
Table VIII presents the FLOPs, number of parameters, and inference time for various methods. Experimental results show that our EME method is the fastest. Interestingly, we notice that a reduction in FLOPs does not directly lead to a similar reduction in the inference time of an approach. This phenomenon may be attributed to the frequent memory access by operators, as confirmed in PConv[68]. This observation inspires us to further speed up our EME method by reducing memory access in the future.
V Conclusions
In this paper, we propose the effective multispectral early-fusion (EME) detector, which achieves both high performance and efficiency. We identify and address performance obstacles in a plain early-fusion strategy, such as information interference, domain gaps, and weak feature representation, by proposing solutions including shape-priority early-fusion modules, weakly supervised learning, and core knowledge distillation. Extensive experiments on representative datasets demonstrate the effectiveness and efficiency of our EME detector.
The main advantage of our EME detector is that it improves the performance of an efficient single-branch early-fusion strategy without significantly increasing its computational burden. We demonstrate that our EME detector has similar FLOPs and parameters to a plain early-fusion strategy, while achieving better performance than a cumbersome two-branch structure. We also show that our EME detector has higher inference efficiency than the two-branch structure on an edge device: the NVIDIA AGX Orin.
A limitation of our current EME detector is its inefficient two-stage training paradigm in the knowledge distillation technique. In the future, we will work towards an optimized one-stage paradigm to accelerate the training process and further improve detection accuracy.